# MH-GIN: Multi-scale Heterogeneous Graph-based Imputation Network for AIS Data (Extended Version)

### v1

@@ Abstract.
doi:XX.XX/XXX.XX Tianyi Li is the corresponding author.
@@ 2. Related Work
Trajectory Imputation. Trajectory imputation methods reconstruct missing segments in the trajectories of moving objects (e.g., vehicles, pedestrians, and ships), typically leveraging motion continuity and spatio-temporal context. Transformer-based approaches, such as TrajBERT (Si et al., 2024), embed geographic coordinates as tokens and reconstruct missing positions using bidirectional attention mechanisms. KAMEL (Musleh and Mokbel, 2023) maps trajectory imputation to the missing word problem in natural language processing, adapting BERT with spatial-awareness and multi-point imputation capabilities for scalable trajectory reconstruction. Recurrent and graph-based sequence models (Xu et al., 2023; Zhao et al., 2022; Yoon et al., 2019; Lipton et al., 2016) exploit temporal dependencies and spatial constraints, capturing complex motion patterns and environmental context to impute missing trajectory segments. Meanwhile, generative models like GANs (Shi et al., 2023) and VAEs (Qi et al., 2020) produce plausible trajectories consistent with known movement behaviors. Specifically tailored maritime approaches include Multi-task AIS (Nguyen et al., 2018), which integrates recurrent neural networks with latent-variable modeling to simultaneously address trajectory reconstruction, anomaly detection, and vessel-type identification. PG-DPM (Zhang et al., 2024) introduces a physics-guided diffusion probabilistic model specifically designed for long-term vessel trajectory imputation by incorporating maritime domain knowledge into the generative process. However, these methods generally focus exclusively on reconstructing vessel positions and rarely consider other heterogeneous attributes (e.g., navigation status or draught), which are critical for comprehensive AIS data analysis.
@@ 3.1. Data and Notation
Spatio-temporal attributes exhibit spatial and temporal characteristics. For instance, spatial scaling varies with latitude: at the equator ($\phi=0^{\circ}$), a longitudinal change of $1^{\circ}$ corresponds to approximately 111 km, while at a latitude of $\phi=79^{\circ}$ North, the same $1^{\circ}$ shift corresponds to roughly 20 km. Moreover, longitude values of $\lambda=180^{\circ}$ and $\lambda=-180^{\circ}$ represent the same geographic meridian, due to spherical continuity. Temporally, timestamps demonstrate nested periodicities — both daily (24×360024$\times$ 360024 × 3600 seconds) and weekly (168×3600168$\times$ 3600168 × 3600 seconds) cycles return to the starting point ($\tau=0$).
@@
Continuous attributes, such as draught ($d$) and speed over ground ($s$), differ substantially in their numerical ranges. Draught typically ranges from 0 to 202020 meters, while speed ranges from 0 to 303030 knots or higher. Due to the distinct scales, a unit change in draught and a unit change in speed are not directly comparable, thus requiring normalization prior to unified analyses.
Discrete attributes represent categorical values without inherent numerical order. For example, the navigation status attribute ($\eta$) includes distinct categories such as 0 (under way), 111 (anchored), and 222 (not under command). These values function purely as categorical labels rather than as numerical measurements.
@@ 3.2. Multi-scale Heterogeneous Graph
[…]
@@ 3.3. Problem Formulation
  • Cross-scale Propagation Phase ($\hat{\mathbf{H}}\leftarrow\mathcal{F}_{3}(\widetilde{\mathbf{H}})$): Propagate representations $\hat{\mathbf{H}}$ across different time scales within attribute subgraphs $\{\mathcal{G}_{x}\}_{x\in\mathcal{X}}$ to impute missing features.
  • Missing Attribute Imputation Phase ($\widetilde{\mathbf{X}}\leftarrow\mathcal{F}_{4}(\mathbf{H},\widetilde{\mathbf{H}},\hat{\mathbf{H}})$): Recover missing attributes $\widetilde{\mathbf{X}}$ based on $\mathbf{H}^{*}$ that combines the original representations $\mathbf{H}$ and propagated representations $\widetilde{\mathbf{H}}$ and $\hat{\mathbf{H}}$ from intra-scale and cross-scale propagation.
@@ 4.1. Heterogeneous Attribute Encoder
Spatio-temporal Attribute Encoder. Spherical coordinates (longitude $\lambda$, latitude $\phi$, as defined in Definition 3.1 and Table 2) exhibit non-Euclidean geometry, as standard numeric encodings fails to capture spherical continuity and latitude-dependent distortion (Cohen et al., 2018). We address this by extending 3D Cartesian projections (Wikipedia contributors, 2024a) with additional harmonic terms to enhance directional sensitivity. We then apply learnable affine transformations with hyperbolic tangent activations to produce embeddings $\mathbf{e}_{\lambda}$ and $\mathbf{e}_{\phi}$.
@@ 4.2. Hierarchical Temporal Feature Extractor
[…]
@@ 4.3. Multi-scale Dependency Mining Module
This property ensures that small perturbations to the input features (e.g., from missing data or noise) result in predictably bounded changes in the final output. The Lipschitz continuity arises from the composition of linear operators with bounded spectral norms, making the model robust to input uncertainties commonly encountered in real-world AIS data. (proof in Appendix B.2 (Liu et al., 2025)).
@@ 4.4. Heterogeneous Attribute Imputation
[…]
@@ 4.5. Training
where $\lambda_{\textit{coo}}$, $\lambda_{\tau}$, $\lambda_{\textit{period}}$, $\lambda_{\textit{cont}}$, and $\lambda_{\textit{disc}}$ are balancing hyperparameters. The training employs specialized loss functions tailored to each attribute type: Haversine distance $\mathcal{L}_{\textit{coo}}$ (Wikipedia contributors, 2024b) for spatio-temporal coordinates to capture spherical distance, mean squared error $\mathcal{L}_{\tau}$ for timestamp intervals, trigonometric encoding loss $\mathcal{L}_{\textit{period}}$ for cyclical attributes to preserve periodicity, mean squared error $\mathcal{L}_{\textit{cont}}$ for continuous attributes, and cross-entropy loss $\mathcal{L}_{\textit{disc}}$ for discrete classification. This multi-objective design ensures that each attribute type is reconstructed according to its inherent characteristics and constraints.
@@ 5.1. Overall Settings
We use two AIS datasets: AIS-DK from the Danish Maritime Authority [^3] http://web.ais.dk/aisdata/ and AIS-US from NOAA [^4] https://coast.noaa.gov/htdata/CMSP/AISDataHandler/2024/index.html. AIS-DK covers Danish waters from March to December 2024, including major shipping routes in the Baltic and North Seas. It contains 18,481 vessel sequences and 10,518,249 AIS records, with an average sequence duration of 9 days. AIS-US spans March to May 2024, focusing on US coastal waters with dense traffic near major ports. It includes 19,091 vessel sequences and 10,546,297 AIS records, with an average sequence duration of 4 days.
@@ A.1. Heterogeneous Attribute Encoder
Spatio-temporal Attribute Encoder. Spherical coordinates (longitude $\lambda$, latitude $\phi$, as defined in Definition 3.1 and Table 2) exhibit non-Euclidean geometry, as standard numeric encodings fails to capture spherical continuity and latitude-dependent distortion (Cohen et al., 2018). We address this by extending 3D Cartesian projections (Wikipedia contributors, 2024a) with additional low-order harmonic terms:
@@
[…]
@@ A.2. Heterogeneous Attribute Imputation
[…]
@@
[…]
@@
[…]
@@ A.3. Training Details
Spatio-temporal Loss. For spatio-temporal attributes, we employ the Haversine distance (Wikipedia contributors, 2024b) to measure the spherical distance between predicted coordinates and ground truth coordinates:
@@
where $(\hat{\lambda},\hat{\phi})$ represents the predicted coordinates, $(\lambda,\phi)$ denotes the ground truth coordinates, and the distance is measured in radians to balance the magnitude of this loss term with other loss components.
@@ B.1. Stability of Multi-scale Propagation
where $\mathbf{P}^{k}=(\mathbf{D}^{k})^{-1/2}\hat{\mathbf{A}}^{k}(\mathbf{D}^{k})^{-1/2}$ is the normalized propagation matrix. The symmetric normalization ensures that $\mathbf{P}^{k}$ is a doubly stochastic matrix when $\hat{\mathbf{A}}^{k}$ represents valid edge weights. By the Perron-Frobenius theorem (Wikipedia contributors, 2025b), the spectral radius of such matrices is bounded: $\rho(\mathbf{P}^{k})\leq 1$. This makes the intra-scale propagation a non-expansive mapping:
@@ B.2. Robustness of Multi-scale Propagation
A function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $L$-Lipschitz continuous if there exists a constant $L\geq 0$ such that for any two inputs $\mathbf{h}_{1},\mathbf{h}_{2}\in\mathbb{R}^{n}$, the inequality $\|g(\mathbf{h}_{1})-g(\mathbf{h}_{2})\|\leq L\|\mathbf{h}_{1}-\mathbf{h}_{2}\|$ holds (Wikipedia contributors, 2025a).
@@ C.1.1. Evaluation Metrics
For spatio-temporal attributes, we implement domain-specific evaluation approaches. For timestamp, we evaluate time intervals between consecutive records using MAE and SMAPE rather than absolute timestamps. For coordinate attributes (latitude and longitude), we compute the spherical distance between true and imputed coordinates in radians using haversine distance (Wikipedia contributors, 2024b).
@@ C.1.3. Masking Strategies
  • Block Masking: Applied to the attributes in time scales 3 and 4 (i.e., $\mathbf{X}^{3}=\{\eta\}$ and $\mathbf{X}^{4}=\{\chi,d\}$). For each voyage segment $[t_{\text{start}},t_{\text{end}}]$ and attribute $x\in\mathbf{X}^{3}\cup\mathbf{X}^{4}$, we randomly mask continuous segments with probability $r$ according to:
@@ C.1.4. Noise Injection Strategies
For timestamp $\tau$, consistent with our model’s method of predicting time intervals (as detailed in Section 4.4 and modeled as a temporal point process (Lin et al., 2022; Omi et al., 2019)), we corrupt the time intervals ($\Delta\tau_{i}$) rather than the absolute timestamps. This directly tests the model’s core mechanism for handling time. Noise is scaled by the standard deviation of these intervals ($\sigma_{\Delta\tau}$) to reflect typical reporting frequency fluctuations. A $\max(0,...)$ function is applied as a necessary physical constraint to prevent time from moving backwards. The noise is then added to the time value $x_{n}$, as follows:
@@
Cyclical Attributes. To model the jitter commonly found in compass and directional sensor readings, we add Gaussian noise to the angular values. Scaling this noise by the standard deviation of the step-to-step angular change ($\sigma_{\Delta x_{c}}$) makes the noise magnitude adaptive to the vessel’s turning dynamics. A modulo 360 operation is applied to ensure that the resulting values remain within the valid [0,360)[0,360)[ 0 , 360 ) range and correctly handle the wrap-around nature of cyclical data. The noise is then added to the angular value, $x_{c}$, as follows:
@@
Discrete Attributes. Errors in categorical data are not numerical deviations but misclassifications (e.g., “anchored” is recorded as “moored”). The most direct way to simulate this is via label flipping. Here, $\gamma$ defines the probability that a given label is flipped to another randomly chosen, valid label. While the role of $\gamma$ here (controlling corruption frequency) differs from its role in other types (controlling corruption magnitude), this preserves $\gamma$ as a unified parameter. It operates on the same [0,1][0,1][ 0 , 1 ] scale, where $\gamma=0$ means no noise and $\gamma=1$ implies maximum corruption for all attribute types. This method allows for a systematic and synchronized increase in domain-appropriate noise across the entire heterogeneous dataset. The noise is then added to the categorical value, $x_{d}$, as follows:

### v3

@@ Abstract.
doi:$X$.XX/XXX.XX Tianyi Li is the corresponding author.
@@ 2. Related Work
Trajectory Imputation. Trajectory imputation methods reconstruct missing segments in the trajectories of moving objects (e.g., vehicles, pedestrians, and ships), typically leveraging motion continuity and spatio-temporal context. Transformer-based approaches, such as TrajBERT (Si et al., 2024), embed geographic coordinates as tokens and reconstruct missing positions using bidirectional attention mechanisms. KAMEL (Musleh and Mokbel, 2023) maps trajectory imputation to the missing word problem in natural language processing, adapting BERT with spatial-awareness and multi-point imputation capabilities for scalable trajectory reconstruction. Recurrent and graph-based sequence models (Xu et al., 2023; Zhao et al., 2022; Yoon et al., 2019; Lipton et al., 2016) exploit temporal dependencies and spatial constraints, capturing complex motion patterns and environmental context to impute missing trajectory segments. Meanwhile, generative models like GANs (Shi et al., 2023) and VAEs (Qi et al., 2020) produce plausible trajectories consistent with known movement behaviors. Specifically tailored maritime approaches include Multi-task AIS (Nguyen et al., 2018), which integrates recurrent neural networks with latent-variable modeling to simultaneously address trajectory reconstruction, anomaly detection, and vessel-type identification. PG-DPM (Zhang et al., 2024) introduces a physics-guided diffusion probabilistic model specifically designed for long-term vessel trajectory imputation by incorporating maritime domain knowledge into the generative process. Complementary studies on trajectory compression and clustering (Li et al., 2020, 2021, 2022; Song et al., 2024) share similar goals of preserving motion patterns and data utility. However, these methods generally focus exclusively on reconstructing vessel positions and rarely consider other heterogeneous attributes, which are critical for comprehensive AIS data analysis.
@@ 3.1. Data and Notation
Spatio-temporal attributes exhibit spatial and temporal characteristics. For instance, spatial scaling varies with latitude: at the equator ($\phi=0^{\circ}$), a longitudinal change of $1^{\circ}$ corresponds to approximately 111 km, while at a latitude of $\phi=79^{\circ}$ North, the same $1^{\circ}$ shift corresponds to roughly 20 km. Moreover, longitude values of $\lambda=180^{\circ}$ and $\lambda=-180^{\circ}$ represent the same geographic meridian, due to spherical continuity. Temporally, timestamps demonstrate nested periodicities — both daily ($24\times 3600$ seconds) and weekly ($168\times 3600$ seconds) cycles return to the starting point ($\tau=0$).
@@
Continuous attributes, such as draught ($d$) and speed over ground ($s$), differ substantially in their numerical ranges. Draught typically ranges from 0 to $20$ meters, while speed ranges from 0 to $30$ knots or higher. Due to the distinct scales, a unit change in draught and a unit change in speed are not directly comparable, thus requiring normalization prior to unified analyses.
Discrete attributes represent categorical values without inherent numerical order. For example, the navigation status attribute ($\eta$) includes distinct categories such as 0 (under way), $1$ (anchored), and $2$ (not under command). These values function purely as categorical labels rather than as numerical measurements.
@@ 3.2. Multi-scale Heterogeneous Graph
AIS data attributes generate hierarchical features across different time scales. Specifically, for an attribute $x^{j}\in\mathbf{X}^{j}$ at its original time scale $j$, features $v_{x^{j}}^{k}\in\mathcal{V}^{k}$ can be derived at time scales $k$, where $k\geq j$. Thus, each attribute $x^{j}$ generates 5+1j5+1-j multi-scale features that capture its temporal features at different time scales.
@@ 3.3. Problem Formulation
[…]
@@ 4.1. Heterogeneous Attribute Encoder
Spatio-temporal Attribute Encoder. Spherical coordinates (longitude $\lambda$, latitude $\phi$, as defined in Definition 3.1 and Table 2) exhibit non-Euclidean geometry, as standard numeric encodings fails to capture spherical continuity and latitude-dependent distortion (Cohen et al., 2018). We address this by extending 3D Cartesian projections (Wikipedia contributors, 2024) with additional harmonic terms to enhance directional sensitivity. We then apply learnable affine transformations with hyperbolic tangent activations to produce embeddings $\mathbf{e}_{\lambda}$ and $\mathbf{e}_{\phi}$.
@@ 4.2. Hierarchical Temporal Feature Extractor
For an attribute $x^{k}$ at time scale $k\in\{1,\ldots,5\}$, we use 5+1k5+1-k recurrent layers, so that higher time-scale attributes require fewer layers. The state update is governed by:
@@ 4.3. Multi-scale Dependency Mining Module
This property ensures that small perturbations to the input features (e.g., from missing data or noise) result in predictably bounded changes in the final output. The Lipschitz continuity arises from the composition of linear operators with bounded spectral norms, making the model robust to input uncertainties commonly encountered in real-world AIS data (proof in Appendix B.2).
@@ 4.4. Heterogeneous Attribute Imputation
Discrete Decoder. For discrete attributes, we apply a linear layer followed by softmax to predict class probabilities $\hat{y}$ from the gated fusion embedding $\mathbf{\tilde{e}}_{x_{d}}$, where the output dimension corresponds to the number of discrete classes $|C|$.
@@ 4.5. Training
where $\lambda_{\textit{coo}}$, $\lambda_{\tau}$, $\lambda_{\textit{period}}$, $\lambda_{\textit{cont}}$, and $\lambda_{\textit{disc}}$ are balancing hyperparameters. The training employs specialized loss functions tailored to each attribute type: Haversine distance $\mathcal{L}_{\textit{coo}}$ (Wikipedia contributors, 2025a) for spatio-temporal coordinates to capture spherical distance, mean squared error $\mathcal{L}_{\tau}$ for timestamp intervals, trigonometric encoding loss $\mathcal{L}_{\textit{period}}$ for cyclical attributes to preserve periodicity, mean squared error $\mathcal{L}_{\textit{cont}}$ for continuous attributes, and cross-entropy loss $\mathcal{L}_{\textit{disc}}$ for discrete classification. This multi-objective design ensures that each attribute type is reconstructed according to its inherent characteristics and constraints.
@@ 5.1. Overall Settings
We use two AIS datasets: AIS-DK from the Danish Maritime Authority [^3] http://aisdata.ais.dk/?prefix=2024/(Last access: 2025.10) and AIS-US from NOAA [^4] https://coast.noaa.gov/htdata/CMSP/AISDataHandler/2024(Last access: 2025.10). AIS-DK covers Danish waters from March to December 2024, including major shipping routes in the Baltic and North Seas. It contains 18,481 vessel sequences and 10,518,249 AIS records, with an average sequence duration of 9 days. AIS-US spans March to May 2024, focusing on US coastal waters with dense traffic near major ports. It includes 19,091 vessel sequences and 10,546,297 AIS records, with an average sequence duration of 4 days.
@@ 6. Conclusion and Future Work
7. Acknowledgment
The research leading to the results presented in this paper has received funding from the European Union’s funded Project MobiSpaces under grant agreement no 101070279.
@@ A.1. Heterogeneous Attribute Encoder
Spatio-temporal Attribute Encoder. Spherical coordinates (longitude $\lambda$, latitude $\phi$, as defined in Definition 3.1 and Table 2) exhibit non-Euclidean geometry, as standard numeric encodings fails to capture spherical continuity and latitude-dependent distortion (Cohen et al., 2018). We address this by extending 3D Cartesian projections (Wikipedia contributors, 2024) with additional low-order harmonic terms:
@@
where $\mathbf{W}_{d}\in\mathbb{R}^{d\times|C|}$, $|C|$ is the number of categories in attribute $x_{d}$, and $x_{d}$ represents any attribute within the discrete attribute set $\mathbf{X}_{d}$. Finally, all type-specific representations are concatenated into:
@@ A.2. Heterogeneous Attribute Imputation
where $\mathbf{W}_{\tau,1},\mathbf{W}_{\tau,2}\in\mathbb{R}^{d\times d}$ are learnable weights, $\mathbf{b}_{\tau,1},\mathbf{b}_{\tau,2}\in\mathbb{R}^{d}$ are bias vectors, and $\text{SiLU}(\cdot)$ is the self-gated activation function (Ramachandran et al., 2018). The predicted timestamp is derived as:
@@
where $\mathbf{W}_{n}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}_{n}\in\mathbb{R}$ are learnable parameters, $\text{ReLU}(\cdot)$ is the Rectified Linear Unit activation function, $\mu,\sigma$ represent the attribute’s statistical moments from training data, and $\alpha,\beta$ are the normalization parameters from the encoder (see Section 4.1).
@@
where $\mathbf{W}_{d}\in\mathbb{R}^{|C|\times d}$ and $\mathbf{b}_{d}\in\mathbb{R}^{|C|}$ are learnable parameters, and $|C|$ is the number of discrete classes.
@@ A.3. Training Details
Spatio-temporal Loss. For spatio-temporal attributes, we employ the Haversine distance (Wikipedia contributors, 2025a) to measure the spherical distance between predicted coordinates and ground truth coordinates:
@@
[…]
@@ B.1. Stability of Multi-scale Propagation
where $\mathbf{P}^{k}=(\mathbf{D}^{k})^{-1/2}\hat{\mathbf{A}}^{k}(\mathbf{D}^{k})^{-1/2}$ is the normalized propagation matrix. The symmetric normalization ensures that $\mathbf{P}^{k}$ is a doubly stochastic matrix when $\hat{\mathbf{A}}^{k}$ represents valid edge weights. By the Perron-Frobenius theorem (Wikipedia contributors, 2025c), the spectral radius of such matrices is bounded: $\rho(\mathbf{P}^{k})\leq 1$. This makes the intra-scale propagation a non-expansive mapping:
@@ B.2. Robustness of Multi-scale Propagation
A function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $L$-Lipschitz continuous if there exists a constant $L\geq 0$ such that for any two inputs $\mathbf{h}_{1},\mathbf{h}_{2}\in\mathbb{R}^{n}$, the inequality $\|g(\mathbf{h}_{1})-g(\mathbf{h}_{2})\|\leq L\|\mathbf{h}_{1}-\mathbf{h}_{2}\|$ holds (Wikipedia contributors, 2025b).
@@ C.1.1. Evaluation Metrics
For spatio-temporal attributes, we implement domain-specific evaluation approaches. For timestamp, we evaluate time intervals between consecutive records using MAE and SMAPE rather than absolute timestamps. For coordinate attributes (latitude and longitude), we compute the spherical distance between true and imputed coordinates in radians using haversine distance (Wikipedia contributors, 2025a).
@@ C.1.3. Masking Strategies
  • Block Masking: Applied to the attributes in time scales 3 and 4 (i.e., $\mathbf{X}^{3}=\{\eta\}$ and $\mathbf{X}^{4}=\{\chi,d\}$). For each voyage segment $t_{\text{start}},t_{\text{end}}]$ and attribute $x\in\mathbf{X}^{3}\cup\mathbf{X}^{4}$, we randomly mask continuous segments with probability $r$ according to:
@@ C.1.4. Noise Injection Strategies
[…]
@@
Cyclical Attributes. To model the jitter commonly found in compass and directional sensor readings, we add Gaussian noise to the angular values. Scaling this noise by the standard deviation of the step-to-step angular change ($\sigma_{\Delta x_{c}}$) makes the noise magnitude adaptive to the vessel’s turning dynamics. A modulo 360 operation is applied to ensure that the resulting values remain within the valid $[0,360)$ range and correctly handle the wrap-around nature of cyclical data. The noise is then added to the angular value, $x_{c}$, as follows:
@@
Discrete Attributes. Errors in categorical data are not numerical deviations but misclassifications (e.g., “anchored” is recorded as “moored”). The most direct way to simulate this is via label flipping. Here, $\gamma$ defines the probability that a given label is flipped to another randomly chosen, valid label. While the role of $\gamma$ here (controlling corruption frequency) differs from its role in other types (controlling corruption magnitude), this preserves $\gamma$ as a unified parameter. It operates on the same $[0,1]$ scale, where $\gamma=0$ means no noise and $\gamma=1$ implies maximum corruption for all attribute types. This method allows for a systematic and synchronized increase in domain-appropriate noise across the entire heterogeneous dataset. The noise is then added to the categorical value, $x_{d}$, as follows: